### v2

# Posets are easily testable

Abstract
Alon and Shapira proved that every monotone class (closed under taking subgraphs) of undirected graphs is strongly testable, that is, under the promise that a given graph is either in the class or $\varepsilon$-far from it, there is a test using a constant number of samples (depending on $\varepsilon$ only) that rejects every graph not in the class with probability at least one half, and always accepts a graph in the class. However, their bound on the number of samples is quite large, since they heavily rely on Szemerédi’s regularity lemma. We study the case of posets and show that every monotone class of posets is easily testable, that is, a polynomial number of samples is sufficient. We achieve this via proving a polynomial removal lemma for posets.
We give a simple classification: for every monotone class of posets there is an $h$ such that the class is indistinguishable (every large enough poset in one class is $\varepsilon$-close to a poset in the other class) from the class of posets free of the chain $C_{h}$. This allows to test every monotone class of posets using $O(\varepsilon^{-1})$ samples. The test has two-sided error, but it is almost complete: the probability to refute a poset in the class is polynomially small in the size of the poset.
The analogous results hold for comparability graphs, too.
Keywords: Property testing, polynomial removal lemma, poset
1 Introduction
The relationship between local and global properties of structures is a central theme in combinatorics and computer science. Since the work of Rubinstein and Sudan [23], testing properties by sampling a small number of elements is an emerging research area. A classical result of this kind is the triangle removal lemma by Ruzsa and Szemerédi [24], usually stated in the form that if a graph $G$ admits at most $\delta|V(G)|^{3}$ triangles then it can be made triangle-free by the removal of at most $\varepsilon|V(G)|^{2}$ edges, where $\delta$ depends only on $\varepsilon$. This can be applied to obtain a combinatorial proof of Roth’s theorem [22] on $3$-term arithmetic progressions, while the hypergraph removal lemma has been used to prove Szemerédi’s theorem. Removal lemmas were proved for abelian groups by Green [16], for linear systems of equations by Král, Serra and Vena [19], for local affine-invariant properties by Bhattacharyya, Fischer, Hatami, Hatami and Lovett [10] and for permutations by Klimošová and Král [18], and by Fox and Wei [12], as well.
We say that a property of digraphs is a set of finite digraphs closed under isomorphism. A digraph $G$ is $\varepsilon$-far from having a property $\Phi$ if any digraph $G^{\prime}$ on the vertex set $V(G)$ that differs by at most $\varepsilon|V(G)|^{2}$ edges from $G$ does not have the property $\Phi$ either. A property $\Phi$ is strongly testable if for every $\varepsilon>0$ there exists an $f(\varepsilon)$ such that if the digraph $G$ is $\varepsilon$-far from having the property $\Phi$ then the induced directed subgraph on $f(\varepsilon)$ vertices chosen uniformly at random does not have the property $\Phi$ with probability at least one half, and it always has the property if $G$ does. Alon and Shapira [4] proved that every monotone property of undirected graphs (that is, closed under the removal of edges and vertices) is strongly testable, see Lovász and Szegedy for an analytic approach [20], while Rödl and Schacht generalized this to hypergraphs [21], see also Austin and Tao [8].
Unfortunately, the dependence on $\varepsilon$ can be quite bad already in the case of undirected graphs: the known upper bounds in the Alon-Shapira theorem are wowzer functions due to the iterated involvement of Szemerédi’s regularity lemma. Following Alon and Fox [7] we call a property easily testable if $f(\varepsilon)$ can be bounded by a polynomial of $\frac{1}{\varepsilon}$, else the property is hard. They showed that both testing perfect graphs and testing comparability graphs are hard. Easily testable properties are quite rare, even triangle-free graphs are hard: Behrend’s construction [9] on sets of integers without $3$-term arithmetic progression leads to a lower bound of magnitude $\varepsilon^{c\log(\varepsilon)}$. Alon proved that $H$-freeness is easily testable in the case of undirected graphs if and only if $H$ is bipartite. For forbidden induced subgraphs Alon and Shapira gave a characterization [5], where there are very few easy cases. Testability is usually hard for hypergraphs studied by Gishboliner and Shapira [13] and ordered graphs investigated by Gishboliner and Tomon [14]. An interesting class of properties easy to test are semialgebraic hypergraphs, see Fox, Pach and Zuk [11]. Surprisingly, $3$-colorability and, in general, "partition problems" turned out to be easily testable, see Goldreich, Goldwasser and Ron [15]. Even a conjecture to draw the borderline between easy and hard properties seems beyond reach.
The goal of this paper is to study testability of posets as special digraphs. By a poset we mean a set equipped with a partial order that is anti-reflexive and transitive. Alon, Ben-Eliezer and Fischer [1] proved that hereditary (closed under induced subgraphs) classes of ordered graphs are strongly testable. This implies the removal lemma for posets and that monotone classes of posets are strongly testable in the following way. To every poset with a linear ordering, we can associate the graph on its base set, where distinct elements $x<y$ are adjacent if $x\prec y$. A graph with a linear ordering is associated with a poset if and only if it has no induced subgraph with two edges on three vertices, where the smallest and largest vertices are not adjacent. An alternative of the application of this general result is to follow the proof of Alon and Shapira [4] using the poset version of Szemerédi’s regularity lemma proved by Hladky, Máthé, Patel and Pikhurko [17].
We show that monotone classes of posets (closed under taking subposets) are easily testable. This is equivalent to the following removal lemma with polynomial bound. The height of a finite poset $P$ is the length of its longest chain, while the width is the size of the largest antichain, denoted by $h(P)$ and $w(P)$, respectively. The chain with $h$ elements is denoted by $C_{h}$. Given two finite posets $P,Q$ a mapping $f:Q\rightarrow P$ is a homomorphism if it is order-preserving, i.e., $f(x)\prec f(y)$ for every $x\prec y$. The probability that a uniform random mapping from $Q$ to $P$ is a homomorphism is denoted by $t(Q,P)$. A poset $P$ is called $Q$-free if it does not contain $Q$ as a (not necessarily induced) subposet.
Theorem 1.1.
[Polynomial removal lemma for posets] Consider an $\varepsilon>0$ and a finite poset $Q$ of height at least two. For every finite poset $P$, if $t(Q,P)=\left(\frac{\varepsilon}{2}\right)^{h(Q)w(Q)^{2}}$ then there exists a $Q$-free (moreover, $C_{h(Q)}$-free) subposet of $P$ obtained by deleting at most $\varepsilon|P|^{2}$ edges.
We use this theorem to show that monotone classes of finite posets are easily testable. We will consider the family of finite posets not in the class. To state our precise result we define the height and width of a set of finite posets $\mathcal{P}$ as
$h(\mathcal{P})=\min_{P\in\mathcal{P}}h(P)\quad\quad\quad w(\mathcal{P})=\min_{\begin{subarray}{c}P\in\mathcal{P}:\\ h(P)=h(\mathcal{P})\end{subarray}}w(P).$
Corollary 1.2.
[Easy testability for monotone classes of posets] Consider a family of finite posets $\mathcal{P}$ and a poset $Q\in\mathcal{P}$ with height $h(\mathcal{P})\geq 2$ and width $w(\mathcal{P})$. For every $\varepsilon>0$ and finite poset $P$, if $t(Q,P)=\left(\frac{\varepsilon}{2}\right)^{h(\mathcal{P})w(\mathcal{P})^{2}}$ then there exists a $\mathcal{P}$-free (moreover, $C_{h(\mathcal{P})}$-free) subposet of $P$ obtained by deleting at most $\varepsilon|P|^{2}$ edges.
Algorithm 1 Basic test for $Q$-free posets
the poset $P$
$P^{\prime}\leftarrow$ subposet on $|Q|$ elements chosen uniformly at random
if  $Q$ is a subposet of $P^{\prime}$ then Reject $P$
else Accept $P$
end if
This test always accepts a $Q$-free poset, and rejects a poset $P$ with probability at least $t(Q,P)$. By Theorem 1.1 it is sufficient to iterate this test $\left(\frac{2}{\varepsilon}\right)^{h(Q)w(Q)^{2}}$ times independently to reject a poset $\varepsilon$-far from being $Q$-free with probability at least one half.
Chains will play an important role in more efficient tests for monotone classes of posets: we give a simple classification of these classes from the testing point of view. Two properties $\Phi_{1}$ and $\Phi_{2}$ of posets are indistinguishable if for every $\varepsilon>0$ and $i=1,2$ there exists $N$ such that for every poset $P$ on at least $N$ elements with property $\Phi_{i}$ there exists a poset $P^{\prime}$ on the same set with property $\Phi_{3-i}$ obtained by changing at most $\varepsilon|P|^{2}$ edges of $P$. Since we are interested in monotone properties we only need to allow deleting edges and not to add them.
Theorem 1.3.
[Indistinguishability] Consider a family of finite posets $\mathcal{P}$, set $h=h(\mathcal{P})\geq 2$ and $w=w(\mathcal{P})$. The class of $\mathcal{P}$-free posets and the class of $C_{h}$-free posets are indistinguishable. Namely, every $C_{h}$-free poset is $\mathcal{P}$-free, and if a poset $P$ is $\mathcal{P}$-free then it has a $C_{h}$-free subposet obtained by the removal of at most $2\left(\frac{h^{2}w^{2}}{|P|}\right)^{\frac{1}{hw^{2}}}|P|^{2}$ edges.
Theorem 1.3 motivates a better understanding of the removal lemma for chains and the testing of $C_{h}$-free posets. First we study the basic test with one-sided error. We can also use this test for $C_{h}$-free posets to test $\mathcal{P}$-free posets, where $h=h(\mathcal{P})$. This test is not complete, but the probability to reject a $\mathcal{P}$-free poset turns out to be negligible, $2\left(\frac{h^{2}w^{2}}{|P|}\right)^{\frac{1}{hw^{2}}}\cdot{\binom{h}{2}}$, where $w=w(\mathcal{P})$, since every copy of $C_{h}$ should contain one of the edges removed in Theorem 1.3. If we iterate the test $\left(\frac{2}{\varepsilon}\right)^{h}$ times independently, then the probability to accept a poset $\varepsilon$-far from being $\mathcal{P}$-free is at most one half. On the other hand, the probability to reject a poset that is $\mathcal{P}$-free is at most $2\left(\frac{h^{2}w^{2}}{|P|}\right)^{\frac{1}{hw^{2}}}{\binom{h}{2}}\left(\frac{2}{\varepsilon}\right)^{h}$, and this is negligible if $\varepsilon,h,w$ are fixed and $|P|$ is large enough.
We can get a more efficient test sampling larger subposets instead of iterating the basic test with a constant number of samples.
Algorithm 2 Subposet test for $C_{h}$-free posets with $s$ samples
the poset $P$
$P^{\prime}\leftarrow$ subposet of $s$ elements chosen uniformly at random
if $C_{h}$ is a subposet of $P^{\prime}$ then Reject $P$
else Accept $P$
end if
It turns out that sampling $s=\left\lceil\frac{4\log(h)+4}{2\varepsilon}\right\rceil$ elements is enough to reject posets $\varepsilon$-far from being $C_{h}$-free with probability at least one half, while we always accept $C_{h}$-free posets.
By Theorem 1.3 this test can also be used for testing $\mathcal{P}$-free posets, where $h(\mathcal{P})=h$: it rejects posets $\varepsilon$-far from $\mathcal{P}$-free with probability at least one half at the price of allowing the error of rejecting a $\mathcal{P}$-free poset with negligible probability.
Theorem 1.4.
[The subposet test] Let $h\geq 2$ be an integer, $\varepsilon>0,c>0$ and $P$ a finite poset. If $P$ is $\varepsilon$-far from being $C_{h}$-free then a random subset of $\left\lceil\frac{4\log(h)+4c+1}{2\varepsilon}\right\rceil$ elements chosen independently and uniformly at random contains a copy of $C_{h}$ with probability at least $1-e^{-c}$.
Observe that being $\varepsilon$-far from every $C_{h}$-free poset guarantees that $\varepsilon$ is small, so the number of samples will be large enough.
Remark 1.5.
Every finite poset $P$ is $\frac{1}{2h-2}$-close to be $C_{h}$-free.
Proof.
[…]
Our bound gives the exact order of magnitude on the necessary number of samples for one-sided testing of $C_{h}$-free posets, see Example 2.7.
The comparability graph $G$ associated with a poset $P$ has vertex set $V(G)=P$ and edge set $E(G)=\{(x,y):x\prec y\text{ or }y\prec x\}$. Alon and Fox proved that it is hard to test if a given graph is a comparability graph [7]. However, under the promise that the input graph is a comparability graph we can test monotone classes. All of our results apply to testing monotone classes of comparability graphs, see Section 4.
In a subsequent work we prove that the exact degree is $(h-1)$ in the polynomial removal lemma for chains (and many other structures). Example  2.2 shows that this is sharp. The proof is too technical for this paper to detail here.
In Section 2 we prove the polynomial removal lemma for chains and Theorem 1.4. Section 3 contains the proofs of Theorem 1.1 and Theorem 1.3. Section 4 discusses our results on comparability graphs.
2 Testing chains
First we prove a removal lemma for chains.
Lemma 2.1.
[Removal lemma for chains] For every $\varepsilon>0$, positive integer $h\geq 2$ and every finite poset $P$, if $t(C_{h},P)<\big{(}\frac{\varepsilon}{2}\big{)}^{h}$ then there exists a $C_{h}$-free subposet of $P$ obtained by the removal of at most $\varepsilon|P|^{2}$ edges of $P$.
Polynomial removal lemmas for directed paths have already been obtained by Alon and Shapira [3], but their bound is $O(\varepsilon^{h^{2}})$. We could use their result to get a removal lemma for chains with worse polynomial bound. However, we improve their bound to degree $h$. This is almost the exact degree, as the following example using the Turán theorem shows.
Example 2.2.
[…]
Claim 2.3.
One has to remove at least $\frac{h-1}{\varepsilon}\binom{\varepsilon|P|/(h-1)}{2}$ edges to get a $C_{h}$-free subposet. This is at least $\frac{\varepsilon}{2h}|P|^{2}$ if $|P|>\frac{h^{2}}{\varepsilon}$.
The inequality $t(C_{h},P)\leq\frac{\varepsilon^{h-1}}{h!}$ holds.
The following algorithm will remove the edges in order to get a $C_{h}$-free poset. We will consider a linear ordering of the poset $P$. We may assume that the set of elements of $P$ is $[|P|]=\{1,2\dots,|P|\}$, and every edge goes from a smaller integer to a larger one.
Algorithm 3 Edge removal using a rank function $r$
$h\in\mathds{Z_{+}}$, $\gamma>0$, poset $P$ on $[|P|]$, where if $x\prec y$ then $x<y$
for $y=1,\dots,|P|$ do
     if  $\quad\exists k:\big{|}\{x:x\prec y,r(x)=k\}\big{|}\geq\gamma|P|$ then
         $r(y)\leftarrow 1+\max\left\{k:\big{|}\{x:x\prec y,r(x)=k\}\big{|}\geq\gamma|P|\right\}$
     else
         $r(y)\leftarrow 1$
     end if
end for
for $x\prec y$ do
     if  $r(x)=r(y)$ then
         $E(P)\leftarrow E(P)\setminus\{(x,y)\}$
     else if $r(y)\geq h$ then
         $E(P)\leftarrow E(P)\setminus\{(x,y)\}$
     end if
end for
$P^{\prime}\leftarrow P$
$P^{\prime}$ on vertex set $[|P|]$, edge set $E(P^{\prime})\subseteq E(P)$
11111222113232344325541111122211323234432554
Figure 1: Example for Algorithm 3 with $h=5,\gamma=\frac{1}{11}$. The Hasse diagram of the poset $P$ is on the left, the ranks are written on the elements, the Hasse diagram of the $P^{\prime}$ is on the right.
Analysis of Algorithm 3:
Claim 2.4.
The following holds.
  1. (1)
    The output $P^{\prime}$ is a poset.
  2. (2)
    The output poset $P^{\prime}$ is $C_{h}$-free.
  3. (3)
    The number of edges $x\prec y$ removed such that $r(x)=r(y)$ is at most $\gamma|P|^{2}$.
  4. (4)
    If the number of elements with rank $r(y)\geq h$ is at most $\gamma|P|$, then the number of edges removed by Algorithm 3 in order to get a $C_{h}$-free poset is at most $2\gamma|P|^{2}$.
Proof.
(1) If $x,y,z$ are distinct elements in $P$ with $(x,y)\in E(P^{\prime}),(y,z)\in E(P^{\prime})$ then $(x,z)\in E(P)$, and $r(z)<h$, $r(x)<r(y)<r(z)$ hence $(x,z)\in E(P^{\prime})$.
(2) Note that $r(x)\leq r(y)$ by the transitivity in posets, hence $r$ is increasing on every chain in $P$. Every edge with $r(x)=r(y)$ have been removed, thus $r$ is strictly increasing on every chain in $P^{\prime}$. The poset $P^{\prime}$ is $C_{h}$-free, since the edges ending at those elements, where $r$ is at least $h$, have been removed.
(3) For every $y$ the number of $x$’s with the same value at $r$ can be at most $\gamma|P|$, else $r(y)$ would be greater than $r(x)$. So the number of such removed edges is at most $\gamma|P|^{2}$.
(4) This is a straightforward consequence of the algorithm and (3). ∎
Proof.
(of Lemma 2.1) We run Algorithm 3 with $h,P$ and $\gamma=\frac{\varepsilon}{2}$.
Claim 2.5.
[…]
Proof.
Observe that there are at least $(\gamma|P|)^{r(x)-1}$ chains on $r(x)$ elements ending at $x$ for every $x$ such that $r$ is strictly increasing on these chains.
[…]
(4) of Claim 2.4 proves the lemma. ∎
Now we use the same rank function (defined in Algorithm 3) to optimize the number of samples to test $C_{h}$-free posets.
Proof.
(of Theorem 1.4) We consider a linear extension $<$ of the ordering $\prec$ of the poset $P$. We might assume that the set of elements of $P$ is $[|P|]=\{1,2\dots,|P|\}$, and the linear ordering $<$ is the ordering of the integers. We define $r:P\mapsto\mathbb{N}$ as in Algorithm 3 with $\gamma=\frac{\varepsilon}{2}$.
Let $X$ be a subset of $\left\lceil\frac{4\log(h)+4c+1}{2\varepsilon}\right\rceil=\left\lceil\frac{\log(h)+c}{\gamma}+\frac{1}{4\gamma}\right\rceil$ elements chosen uniformly at random from $P$. We prove that with probability at least $(1-e^{-c})$ there is a chain with elements $x_{h}\prec\dots\prec x_{2}\prec x_{1}$ such that $r(x_{k})=h-k+1$ for all $k\in[h]$. We will find these elements one by one starting with $x_{1}$.
We show that there are at least $\gamma|P|$ elements $x\in P$ with $r(x)=h$. Suppose for a contradiction that there are less. Then running Algorithm 3 gives a $C_{h}$-free poset and by (4) of Claim 3 we deleted at most $\varepsilon|P|^{2}$ edges, contradicting that $P$ was $\varepsilon$-far from being $C_{h}$-free.
Thus, the probability that we do not choose any element with $r(x)=h$ into the set $X$ is at most $(1-\gamma)^{\gamma^{-1}(\log(h)+c+1/4)}<\frac{e^{-c}}{h}$. Denote by $x_{1}$ the smallest element (in the linear extension) such that $r(x_{1})=h$, if there is such an element.
Claim 2.6.
Consider ${x_{1}},{x_{2}},\dots,{x_{k}}$ for $k<h$ such that for every $\ell\in\{2,3,\dots,k\}$ the element $x_{\ell}$ is the smallest (in the linear extension) such that $r(x_{\ell})=h-\ell+1$ and $x_{\ell}\prec x_{\ell-1}$. Then the conditional distribution on the choice of $x_{1},\dots,x_{k}$ of the other elements of $X$ is uniform on the set
$S_{k}:=\left\{x\in P\setminus\{x_{1},x_{2},\dots,x_{k}\}:\forall\ell\in[k]\text{ if }x<x_{\ell}\text{ then either }r(x)\neq h-\ell+1\text{ or }x\nprec x_{\ell-1}\right\}.$
Proof.
Note that $X\subseteq S_{k}\cup\{x_{1},x_{2},\dots,x_{k}\}$: else for the smallest $x$ (in the linear extension) such that $x\notin S_{k}\cup\{x_{1},x_{2},\dots,x_{k}\}$ there would be an $\ell$ such that $x<x_{\ell}$, $r(x)=h-\ell+1$ and $x\prec x_{{\ell-1}}$. Hence we should have chosen $x$ instead of $x_{\ell}$.
On the other hand, the set $X$ could be $S^{\prime}\cup\{x_{1},x_{2},\dots,x_{k}\}$ for any subset $S^{\prime}\subseteq S_{k}$ of size $\left\lceil\frac{\log(h)+c}{\gamma}+\frac{1}{4\gamma}\right\rceil-k$. Since the conditional distribution of $X$ is uniform on these sets the claim follows. ∎
Now we show that a suitable $x_{k+1}$ exists with probability at least $1-\frac{e^{-c}}{h}$.
There are at least $\gamma|P|$ elements $x\in P$ (in particular, $x\in P\setminus\{x_{1},x_{2},\dots,x_{k}\}$ by the partial ordering) such that $x\prec x_{k}$ and $r(x)=h-k$ by the definition of the rank function. Let us denote these good candidates for $x_{k+1}$ by $R_{k+1}$.
Since $\varepsilon<\frac{1}{2h-2}$ and $\gamma<\frac{1}{h-1}$, there are at least $\frac{\log(h)+c}{\gamma}$ elements in $X\setminus\{x_{1},x_{2},\dots,x_{k}\}$. The probability that none of them is in $R_{k+1}$ is at most $(1-\gamma)^{\gamma^{-1}(\log(h)+c)}<\frac{e^{-c}}{h}$. Let $x_{k+1}$ be the smallest element (in the linear extension) such that $x_{k+1}\in R_{k+1}\cap X$, if there is such an element.
The union bound yields the theorem. ∎
The next example shows that Theorem 1.4 gives the exact order of magnitude on the number of samples required for two-sided testing.
The complete $h$-partite poset with antichains of size $w_{1},w_{2},\dots,w_{k}$, respectively, will be denoted by $K_{w_{1},w_{2},\dots,w_{k}}$, while $K_{h\times w}$ is the shorthand notation for $K_{w,w,\dots,w}$ with $hw$ elements. In particular, $K_{h\times 1}$ is the chain $C_{h}$.
Example 2.7.
Given $\varepsilon>0$ and the positive integers $h\geq 2,w\geq 1$ such that $\varepsilon w$ is also an integer, consider the poset $P=K_{\varepsilon w,w,w,\dots,w}$ with $(\varepsilon+h-1)w$ elements.
Claim 2.8.
A subposet obtained by the removal of less than $\varepsilon w^{2}$ edges from $P$ contains $C_{h}$ as a subposet. If we remove all of the $\varepsilon w^{2}$ edges between the first two antichains, we obtain a $C_{h}$-free poset.
Proof.
Every edge with an endvertex in the first antichain (of size $\varepsilon w$) is contained by exactly $w^{h-2}$ chains of height $h$, since we can choose the other elements of the chain from the other antichains arbitrarily. On the other hand, an edge not adjacent to the first antichain is contained by $\varepsilon w^{h-2}$ chains. Since $P$ contains $\varepsilon w^{h}$ chains of height $h$, we need at least $\varepsilon w^{2}$ edges to cover these. ∎
Claim 2.9.
The probability that a random subset with at most $\varepsilon^{-1}$ elements does not contain $C_{h}$ as a subposet is at least $\frac{1}{e}$.
Proof.
Every subposet isomorphic to $C_{h}$ has an element in the first antichain. The probability that the subposet contains no element of this antichain is $\big{(}\frac{h-1}{h-1+\varepsilon}\big{)}^{\varepsilon^{-1}}\geq\big{(}\frac{1}{1+\varepsilon}\big{)}^{\varepsilon^{-1}}>\frac{1}{e}$. ∎
3 Testing monotone classes of posets
The next lemma provides a lower bound on the density of the complete $h$-partite poset $K_{h\times w}$ in terms of the density of the chain of length $h$. The proof uses standard techniques appearing in the solution to the Zarankiewicz problem. We use again the notation $[n]=\{1,2,\dots,n\}$.
Lemma 3.1.
For every poset $P$ and positive integers $h,w$ the inequality
$t(K_{h\times w},P)\geq t^{w^{2}}(C_{h},P)$
holds.
Proof.
The following two claims imply the lemma.
Claim 3.2.
$t(K_{w,1,w,1,\dots},P)\geq t^{w}(C_{h},P)$
Proof.
Note that $K_{w,1,w,1,\dots}$ is the union of $w$ chains of length $h$ intersecting only on the even layers (where it has only one element), and a mapping of $K_{w,1,w,1,\dots}$ is a homomorphism if and only if its restriction to every chain is a homomorphism. Consider a mapping of the even layers of $K_{w,1,w,1,\dots}$. The events that the random mapping gives a homomorphism for chains are conditionally independent for disjoint chains (conditioning on the the mapping of the even layers). Hence the conditional probability that mapping $w$ elements for every odd layer gives a homomorphism of $K_{w,1,w,1,\dots}$ is the $w^{th}$ power of the probability that mapping only one element for every odd layer gives a homomorphism of the chain $C_{h}$. We use Jensen’s inequality to obtain the required result. Now we describe this argument more formally.
Let $\left(x_{i,j}\right)_{i\in[h],j\in[w]\text{ for }i\text{ odd}}$ and $\left(x_{i,1}\right)_{i\in[h]\text{ for }i\text{ even}}$ be chosen uniformly and independently at random.
$t&(K_{w,1,w,1,\dots},P)=\mathds{P}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\text{ for }i\text{ even }\\ \left(x_{i,j}\right)_{i\in[h],j\in[w]}\text{ for }i\text{ odd }\end{subarray}}\left(\Bigg{.}\forall i^{\prime}\in[h-1],j^{\prime}\in[w]\quad\begin{subarray}{c}\text{ if }i^{\prime}\text{ odd then }x_{i^{\prime},j^{\prime}}\prec x_{i^{\prime}+1,1}\\ \text{ if }i^{\prime}\text{ even then }x_{i^{\prime},1}\prec x_{i^{\prime}+1,j^{\prime}}\end{subarray}\right)\\ &=\mathds{E}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\\ i\text{ even }\end{subarray}}\left[\mathds{P}_{\begin{subarray}{c}\left(x_{i,j}\right)_{i\in[h],j\in[w]}\\ i\text{ odd}\end{subarray}}\left(\Bigg{.}\forall i^{\prime}\in[h-1],j^{\prime}\in[w]\quad\begin{subarray}{c}\text{ if }i^{\prime}\text{ odd then }x_{i^{\prime},j^{\prime}}\prec x_{i^{\prime}+1,1}\\ \text{ if }i^{\prime}\text{ even then }x_{i^{\prime},1}\prec x_{i^{\prime}+1,j^{\prime}}\end{subarray}\Big{|}\left(x_{i,1}\right)_{i\in[h]},i\text{ even }\right)\right]\end{aligned}$
Here we split $K_{w,1,w,1,\dots}$ into $w$ edge-disjoint copies of $C_{h}$. Since the events corresponding to elements in the same odd layer are independent we obtain that this equals
$\begin{split}&\mathds{E}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\\ i\text{ even }\end{subarray}}\left[\mathds{P}^{w}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\\ i\text{ odd}\end{subarray}}\left(\Bigg{.}\forall i^{\prime}\in[h-1]\quad x_{i^{\prime},1}\prec x_{i^{\prime}+1,1}\Big{|}\left(x_{i,1}\right)_{i\in[h]},i\text{ even }\right)\right]\\ &\geq\mathds{E}^{w}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\\ i\text{ even }\end{subarray}}\left[\mathds{P}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\\ i\text{ odd}\end{subarray}}\left(\Bigg{.}\forall i^{\prime}\in[h-1]\quad x_{i^{\prime},1}\prec x_{i^{\prime}+1,1}\Big{|}\left(x_{i,1}\right)_{i\in[h]},i\text{ even }\right)\right]\\ &=\mathds{P}^{w}_{\left(x_{i,1}\right)_{i\in[h]}}\left(\Bigg{.}\forall i^{\prime}\in[h-1]\quad x_{i^{\prime},1}\prec x_{i^{\prime}+1,1}\right)=t^{w}(C_{h},P),\end{split}$
where we have applied Jensen’s inequality. ∎
Claim 3.3.
$t(K_{h\times w},P)\geq t^{w}(K_{w,1,w,1,\dots},P)$
Proof.
The proof is very similar to the previous one. Now we use the observation that $K_{h\times w}$ is the union of $w$ copies of $K_{w,1,w,1,\dots}$ intersecting only on the odd layers (where $K_{w,1,w,1,\dots}$ has $w$ elements), and a mapping of $K_{h\times w}$ is a homomorphism if and only if its restriction to every such copy of $K_{w,1,w,1,\dots}$ is a homomorphism. Consider a mapping of the odd layers of $K_{h\times w}$. The events that the random mapping gives a homomorphism for copies of $K_{w,1,w,1,\dots}$ are conditionally independent for disjoint copies of $K_{w,1,w,1,\dots}$ (conditioning on the mapping of the odd layers). Hence the conditional probability that mapping $w$ elements for every even layer gives a homomorphism of $K_{h\times w}$ is the $w^{th}$ power of the probability that mapping only one element for every even layer gives a homomorphism of $K_{w,1,w,1,\dots}$. We use again Jensen’s inequality to obtain the required result.
Let $\left(x_{i,j}\right)_{i\in[h],j\in[w]}$ be chosen uniformly and independently at random.
$t&(K_{h\times w},P)=\mathds{P}_{\left(x_{i,j}\right)_{i\in[h],j\in[w]}}\Big{(}\forall i^{\prime}\in[h-1],j,j^{\prime}\in[w]\quad x_{i^{\prime},j}\prec x_{i^{\prime}+1,j^{\prime}}\Big{)}\\ &=\mathds{E}_{\begin{subarray}{c}\left(x_{i,j}\right)_{i\in[h],j\in[w]}\\ i\text{ odd }\end{subarray}}\left[\mathds{P}_{\begin{subarray}{c}\left(x_{i,j}\right)_{i\in[h],j\in[w]}\\ i\text{ even}\end{subarray}}\left(\forall i^{\prime}\in[h-1],j,j^{\prime}\in[w]\quad x_{i^{\prime},j}\prec x_{i^{\prime}+1,j^{\prime}}\Big{|}\left(x_{i,j}\right)_{i\in[h],j\in[w]},i\text{ odd }\right)\right].\end{aligned}$
Here we split $K_{h\times w}$ into $w$ edge-disjoint copies of $K_{w,1,w,1,\dots}$. Since the events corresponding to elements in the same even layer are independent we obtain that this equals
$&\mathds{E}_{\begin{subarray}{c}\left(x_{i,j}\right)_{i\in[h],j\in[w]}\\ i\text{ odd }\end{subarray}}\left[\mathds{P}^{w}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\\ i\text{ even}\end{subarray}}\left(\Bigg{.}\forall i^{\prime}\in[h-1],j^{\prime}\in[w]\quad\begin{subarray}{c}\text{ if }i^{\prime}\text{ odd then }x_{i^{\prime},j^{\prime}}\prec x_{i^{\prime}+1,1}\\ \text{ if }i^{\prime}\text{ even then }x_{i^{\prime},1}\prec x_{i^{\prime}+1,j^{\prime}}\end{subarray}\Big{|}\left(x_{i,j}\right)_{i\in[h],j\in[w]},i\text{ odd }\right)\right]\\ &\geq\mathds{E}^{w}_{\begin{subarray}{c}\left(x_{i,j}\right)_{i\in[h],j\in[w]}\\ i\text{ odd }\end{subarray}}\left[\mathds{P}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\\ i\text{ even}\end{subarray}}\left(\Bigg{.}\forall i^{\prime}\in[h-1],j^{\prime}\in[w]\quad\begin{subarray}{c}\text{ if }i^{\prime}\text{ odd then }x_{i^{\prime},j^{\prime}}\prec x_{i^{\prime}+1,1}\\ \text{ if }i^{\prime}\text{ even then }x_{i^{\prime},1}\prec x_{i^{\prime}+1,j^{\prime}}\end{subarray}\Big{|}\left(x_{i,j}\right)_{i\in[h],j\in[w]},i\text{ odd }\right)\right]\\ &=\mathds{P}^{w}_{\begin{subarray}{c}\left(x_{i,1}\right)_{i\in[h]}\text{ for }i\text{ even }\\ \left(x_{i,j}\right)_{i\in[h],j\in[w]}\text{ for }i\text{ odd }\end{subarray}}\left(\Bigg{.}\forall i^{\prime}\in[h-1],j^{\prime}\in[w]\quad\begin{subarray}{c}\text{ if }i^{\prime}\text{ odd then }x_{i^{\prime},j^{\prime}}\prec x_{i^{\prime}+1,1}\\ \text{ if }i^{\prime}\text{ even then }x_{i^{\prime},1}\prec x_{i^{\prime}+1,j^{\prime}}\end{subarray}\right)\\ &=t^{w}(K_{w,1,w,1,\dots},P),\end{aligned}$
where we have applied Jensen’s inequality. ∎
The lemma follows. ∎
Proof.
(of Theorem 1.1) Assume that $t(Q,P)<\left(\frac{\varepsilon}{2}\right)^{hw^{2}}$. The poset $Q$ is a subposet of $K_{h\times w}$, so Lemma 3.1 gives $t^{w^{2}}(C_{h},P)\leq t(K_{h\times w},P)\leq t(Q,P)$. These yield $t(C_{h},P)<\left(\frac{\varepsilon}{2}\right)^{h}$, so by Lemma 2.1 there is a $C_{h}$-free subposet $P^{\prime}$ of $P$ obtained by deleting at most $\varepsilon|P|^{2}$ edges. ∎
Proof.
(of Theorem 1.3) If a poset is $C_{h}$-free then it is $\mathcal{P}$-free.
In order to prove the other direction, consider a poset $Q\in\mathcal{P}$ with minimal height $h=h(\mathcal{P})$ and (amongst these) minimal weight $w=w(\mathcal{P})$. If a poset $P$ is $Q$-free then there is no injective homomorphism from $Q$ to $P$, hence $t(Q,P)\leq|P|^{-1}|Q|^{2}$. Since $|Q|\leq hw$, Theorem 1.1 shows that one can get a $C_{h}$-free subposet of $P$ by the removal of $2(h^{2}w^{2})^{\frac{1}{hw^{2}}}|P|^{-\frac{1}{hw^{2}}}|P|^{2}$ edges. ∎
4 Comparability graphs
We will obtain the same theorems for monotone classes of comparability graphs as for posets: the difference will only be in the hidden constants. These allow the same tests as for posets. For a fixed finite graph $F$, the basic test samples $|V(F)|$ vertices and accepts a graph if these do not span an isomorphic copy of $F$. The next removal lemma shows how many iterations we need in order to reject comparability graphs $\varepsilon$-far from being $F$-free with probability at least one half, while always accepting $F$-free comparability graphs. Similarly to posets, given two finite graphs $F,G$ the probability that a uniform random mapping from $F$ to $G$ is a homomorphism (i.e., edge-preserving) is denoted by $t(F,G)$.
Theorem 4.1.
[Polynomial removal lemma for comparability graphs] Consider an $\varepsilon>0$ and a finite graph $F$ that is not an independent set. For every finite comparability graph $G$, if $t(F,G)\leq\left(\frac{\varepsilon}{2}\right)^{\chi(F)\alpha(F)^{2}}$ then there exists an $F$-free (moreover, $K_{\chi(F)}$-free) spanning subgraph of $G$ that is a comparability graph, obtained by deleting at most $\varepsilon|V(G)|^{2}$ edges.
Proof.
The graph $F$ is a subgraph of the multipartite Turán graph $T$ with $\chi(F)$ classes each of size $\alpha(F)$, hence $t(F,G)\geq t(T,G)$. Let $P$ be one of the posets whose comparability graph is $G$. The height of the poset $P$ is exactly the chromatic number of $G$, and the width of the poset $P$ equals the independence number of $G$.
Note that $t(T,G)\geq t(K_{\chi(F)\times\alpha(F)},P)$, since we may assume that $T$ is the comparability graph of $K_{\chi(F)\times\alpha(F)}$, hence every homomorphism of $K_{\chi(F)\times\alpha(F)}$ to $P$ is a comparability-preserving map from $T$ to $G$, i.e., a graph homomorphism. By Theorem 1.1 there exists a $C_{\chi}$-free subposet of $P$ obtained by deleting at most $\varepsilon|P|^{2}$ edges, and its comparability graph satisfies the conditions of the theorem. ∎
Given a set of (possibly infinitely many) finite graphs $\mathcal{F}$ we define the chromatic number $\chi(\mathcal{F})$ and the independence number $\alpha(\mathcal{F})$ as follows.
$\chi(\mathcal{F})=\min_{F\in\mathcal{F}}\chi(F)\quad\quad\quad\alpha(\mathcal{F})=\min_{\begin{subarray}{c}F\in\mathcal{F}:\\ \chi(F)=\chi(\mathcal{F})\end{subarray}}\alpha(F).$
Corollary 4.2.
[Easy testability for monotone classes of comparability graphs] Consider a family of finite graphs $\mathcal{F}$ and a graph $F\in\mathcal{F}$ with chromatic number $\chi(\mathcal{F})\geq 2$ and independence number $\alpha(\mathcal{F})$. For every $\varepsilon>0$ and finite comparability graph $G$, if $t(F,G)\leq\left(\frac{\varepsilon}{2}\right)^{h(\mathcal{F})w(\mathcal{F})^{2}}$ then there exists an $\mathcal{F}$-free (moreover, $K_{\chi(\mathcal{F})}$-free) spanning subgraph of G, that is a comparability graph, obtained by deleting at most $\varepsilon|V(G)|^{2}$ edges.
We also give a classification of monotone classes of comparability graphs as we did for posets. Two properties $\Phi_{1}$ and $\Phi_{2}$ of graphs are indistinguishable if for every $\varepsilon>0$ and $i=1,2$ there exists $N$ such that for every graph $G$ on at least $N$ vertices with property $\Phi_{i}$ there exists a graph $G^{\prime}$ on the same vertex set with property $\Phi_{3-i}$, obtained by changing at most $\varepsilon|V(G)|^{2}$ edges of $G$. Since we are interested in monotone properties we only need to allow deleting edges.
Theorem 4.3.
[Indistinguishability] Consider a family of finite graphs $\mathcal{F}$. Set $\chi=\chi(\mathcal{F})\geq 2,\alpha=\alpha(\mathcal{F})$. Comparability graphs with chromatic number at most $(\chi-1)$ are indistinguishable from $\mathcal{F}$-free comparability graphs. Namely, every comparability graph with chromatic number at most $(\chi-1)$ is $\mathcal{F}$-free, and every $\mathcal{F}$-free comparability graph admits a spanning subgraph with chromatic number at most $(\chi-1)$, that is a comparability graph, obtained by the removal of at most $2\left(\frac{\chi^{2}\alpha^{2}}{|V(G)|}\right)^{\frac{1}{\chi\alpha^{2}}}|V(G)|^{2}$ edges.
Proof.
Clearly every comparability graph with chromatic number at most $(\chi-1)$ is $\mathcal{F}$-free. On the other hand, given an $\mathcal{F}$-free comparability graph $G$ consider a poset $P$ whose comparability graph is $G$. Theorem 1.3 implies that there is a $C_{\chi}$-free subposet $P^{\prime}$ obtained by the removal of at most $2\left(\frac{\chi^{2}\alpha^{2}}{|V(G)|}\right)^{\frac{1}{\chi\alpha^{2}}}|V(G)|^{2}$ edges. The comparability graph $G^{\prime}$ of $P^{\prime}$ is the desired spanning subgraph of $G$: it is $K_{\chi}$-free, since $P^{\prime}$ is $C_{\chi}$-free. Hence $\chi(G^{\prime})\leq\chi-1$ by the dual of the Dilworth theorem. ∎
The analogue of Algorithm 2 is the test sampling a random set of vertices and accepting the graph if the subgraph spanned by them is $K_{\chi}$-free. We need the same number of samples as in the case of posets. The next theorem is a straightforward consequence of Theorem 1.4.
Theorem 4.4.
[On the subgraph test] Let $\chi\geq 2$ be an integer, $\varepsilon>0,c>0$ and $G$ a finite comparability graph. If $G$ is $\varepsilon$-far from being $K_{\chi}$-free then a random subset of $\left\lceil\frac{4\log(\chi)+4c+1}{2\varepsilon}\right\rceil$ vertices chosen independently and uniformly at random contains a copy of $K_{\chi}$ with probability at least $1-e^{-c}$.
The comparability graph of the poset in Example 2.7 shows that this bound has the exact degree. As in the case of posets, we can also use the test for $K_{\chi(\mathcal{F})}$-free subgraphs to test a monotone class of comparability graphs $\mathcal{F}$: the probability that we reject an $\mathcal{F}$-free comparability graph is negligible.
5 Acknowledgements
Panna Tímea Fekete’s Project No. 1016492. has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the KDP-2020 funding scheme and supported by the ERC Synergy Grant No. 810115 – DYNASNET. Gábor Kun was supported by the Hungarian Academy of Sciences Momentum Grant no. 2022-58 and ERC Advanced Grant ERMiD.